### v1

@@ Abstract
Many optimization problems require hyperparameters, i.e., parameters that must be pre-specified in advance, such as regularization parameters and parametric regularizers in variational regularization methods for inverse problems, and dictionaries in compressed sensing. A data-driven approach to determine appropriate hyperparameter values is via a nested optimization framework known as bilevel learning. Even when it is possible to employ a gradient-based solver to the bilevel optimization problem, construction of the gradients, known as hypergradients, is computationally challenging, each one requiring both a solution of a minimization problem and a linear system solve. These systems do not change much during the iterations, which motivates us to apply recycling Krylov subspace methods, wherein information from one linear system solve is re-used to solve the next linear system. Existing recycling strategies often employ eigenvector approximations called Ritz vectors. In this work we propose a novel recycling strategy based on a new concept, Ritz generalized singular vectors, which acknowledge the bilevel setting. Additionally, while existing iterative methods primarily terminate according to the residual norm, this new concept allows us to define a new stopping criterion that directly approximates the error of the associated hypergradient. The proposed approach is validated through extensive numerical testing in the context of an inverse problem in imaging.
@@ 1 Introduction
Large-scale optimization problems that arise in applications such as machine learning [19, 37, 61, 63] and inverse problems [4, 17, 26, 55] often require the pre-specification of certain parameters: dictionaries for compressed sensing [11, 36, 40], sample weights for a biased dataset [50, 56]. Referred to as hyperparameters, it is crucial that appropriate values of these parameters are determined in order to ensure both convergence of the optimization process and that a meaningful solution is obtained. For example, for inverse problems, which arise in many important science and engineering applications such as biomedical, astronomical, and seismic imaging [3, 4, 17, 27], one is tasked to recover some quantity of interest given and indirect measurements thereof. In practice, a common approach to determine a meaningful solution is to solve an optimization problem [4, 17, 26, 55] such as
where the regularizer is used to encourage a-priori information of the solution in the reconstruction, such as sparsity in some signal domain. Here the regularizer is parametrized by some vectors which may represent, for example: the weights for a sum of different regularizers [12]; convolution filters [7, 51]; parameters of a neural network [2, 24, 25, 38]. Determining appropriate values of is crucial to attain a meaningful solution, as a poor choice can lead to a reconstruction that is dominated by noise or over-smoothed [26].
@@
i.e., a linear system whose coefficient matrix is the lower level Hessian. Thus, a gradient-based method applied to (2) involves the solution of a sequence of linear systems of the form (3), which stabilize as the iterations for solving (2) proceed. In applications such as inverse problems, each linear system (3) is typically large scale and thus solved by an iterative method. Beyond warm-starting, a way to exploit similarity in a sequence of linear systems is to employ recycling Krylov subspace methods [46, 59, 62], a class of iterative solvers wherein the solution space of one linear system is ‘recycled’ in the solve of the next linear system.
The remaining part of this paper is organized as follows. In Section 2 we provide some background material on gradient-based methods for the solution of bilevel optimization problems, as well as recycling Krylov methods for a sequence of linear systems. In Section 3 we first briefly summarize existing strategies to determine the subspace to be recycled, which are primarily based on Ritz vectors, and then introduce Ritz generalized singular vectors that lead to a novel recycling strategy acknowledging that, in the bilevel optimization setting, solutions of the linear systems are employed in the construction of hypergradients. In Section 4 we show how our proposed recycling strategy allows for a stopping criterion based on approximations of the hypergradient error, as opposed to existing stopping criteria based on the Hessian system residual in some norm [20, 49, 62]. Section 5 reports the results of extensive numerical experiments performed on a test problem, where we show the benefit of using recycling Krylov methods, and compare existing and our new recycling strategy and stopping criterion. Section 6 presents some concluding remarks.
@@ 2.1 Gradient-based solvers for bilevel optimization
where is the Jacobian of with respect to We typically select an upper level loss function such that is easy to compute, e.g., yields An issue with (5) is the application of the Jacobian Indeed, even with existence and uniqueness of we often lack an analytic solution for which we can compute the Jacobian of. However, by the assumptions on an alternative form of the Jacobian motivated by the implicit function theorem [10, 13, 16, 66] can be employed. Namely,
@@
Regarding the computational feasibility of (7), for applications such as imaging problems (whereby is a vectorialization of inherently multi-dimensional quantities) the Hessian is very large and, in general, lacks any structure. Indeed, we may only have access to a function that implicitly applies the action of the Hessian onto a vector. For these reasons the Hessian is not explicitly inverted in practice, and instead an iterative method is employed to compute the solution of the linear system
@@
At iteration of a gradient-based solver for (4), having the current solution and hypergradient (7) (acting as a search direction ), an appropriate steplength can be determined via backtracking linesearch with the Armijo stopping rule. Namely, we find the smallest such that
@@
The caveat of applying the solver described in the previous subsection is that, in order to form each hypergradient (9), one must determine the solution to both a sequence of lower level problems (4b) and Hessian system (8). However, as the iterations of the gradient-based optimization scheme proceed, one would expect the choice of hyperparameters to stabilise; in particular the associated sequences of solutions to the lower level and Hessian systems would also stabilise.
We focus our attention on the sequence of linear systems that are solved. To compact notation with respect to (8), we denote the Hessian system encountered at the th iteration of a gradient-based method for solving (4), that is, line 4 of Algorithm (2), as
@@ 2.2.1 MINRES
We now focus on how a given linear system in (11) can be solved. To further compact notation, we suppress the superscript in (11) and consider the linear system with initial guess One way to solve the linear system is via a Krylov subspace method [22, 52], such as MINRES [44], which builds a Krylov subspace where is the initial residual, and computes a solution that is optimal over
@@
where is the first canonical basis vector. If the residual norm is sufficiently small the algorithm can be terminated, otherwise a new basis vector can be efficiently constructed by the 3-term recurrence (12) and the associated residual norm determined. In particular, MINRES reduces the solution of the original linear system to the solution of a tridiagonal system of order , which can be efficiently performed; indeed, the Lanczos relation (12) allows the QR decomposition of to be efficiently updated to yield the QR decomposition of ; see [44] for details.
@@ 2.2.2 RMINRES
A recycling variant of MINRES, RMINRES [62], finds a correction term of the form
@@
where is a pre-specified recycle space and the columns of form an orthonormal basis of a Krylov subspace which is iteratively constructed. We remind ourselves that, although we suppress the notation in this subsection, every variable in (14) - including the recycle space - depends on the iteration of the gradient-based algorithm employed to solve the bilevel problem.
@@
where can be shown to be tridiagonal [59, Section 5.2]. This precise choice of orthogonalisation allows (15) to be decoupled into
@@ 3.1 Existing selection criteria
Existing works that employ recycling strategies are primarily motivated by two settings: a sequence of slowly changing linear systems are solved; or a single linear system is solved but, due to computational memory restrictions, the numerical solver must be restarted mid-solve [41]. As we are interested in solving (11), we tailor discussion to the former. With this in mind, we interpret as the Hessian of the system we are about to solve, we include a discussion on the implications of this choice later in the section.
Although the two settings are different, existing recycling strategies are motivated by approximating the eigendecomposition
where is orthonormal and is a diagonal matrix of the eigenvalues. Indeed, if one has access to the full eigendecomposition (19) then the correction term is Due to the large-scale nature of the system and, in general, lack of structure of the Hessian, it is not computationally feasible to compute the eigendecomposition for every Hessian system, hence an iterative method is employed. Even so, since is in the column space of choosing to mimic in some capacity is a natural choice.
@@
and consider a subspace in the range of The columns of are called Ritz vectors, and the associated eigenvalues (entires of the diagonal matrix ) are called the Ritz values. Notice that should we recover the full eigendecomposition Ritz values tend to approximate the extremal eigenvalues of well, but can give poor approximations to interior eigenvalues [42]. For example, eigenvalues can be written as a min-max characterisation of the Rayleigh quotient [47, Theorem 10.2.1] and one of the Ritz values is the largest possible Rayleigh quotient over the given subspace [47, Theorem 11.4.1]. Other characterizations of how Ritz vectors/values can be considered optimal approximations of eigenvectors/values over the subspace spanned by are given in [23, Chapter 8].
It turns out that the Lanczos algorithm for generating Krylov basis vectors is the Rayleigh-Ritz procedure applied to a Krylov subspace [47, Chapter 13], and so the generated Krylov subspace inherently approximates exterior eigenvectors. For this reason, recycling approximations of interior eigenvectors can propagate information least readily available from the Lanczos algorithm. An approximation of interior eigenvalues are Harmonic Ritz values, which are the reciprocal of (ordinary) Ritz values of with respect to Harmonic Ritz values are named as such due to the interpretation of them as a weighted harmonic mean of ’s eigenvalues, in contrast to Ritz values which can be viewed as a weighted arithmetic mean of ’s eigenvalues; see [43] for details. In practice, harmonic Ritz vectors are not determined by forming and computing the associated Ritz vectors, but instead an equivalent generalized eigenvalue problem
@@
Other recycling strategies are possible, for example, in [33] the recycling subspace is selected by choosing the columns of that correspond to the coefficients in of largest magnitude, as it is an indication of how important a given column of is in representing the solution.
@@ 3.2 Acknowledging the bilevel setting
where In practice we first approximate and then perform a premultiplication by The recycling strategies outlined in Section 3.1 use vectors related to a projected approximation of and so are oblivious to such a premultiplication to the determined solution. To acknowledge the bilevel setting, we propose to approximate over the column-space of a full-rank matrix by using the Generalized Singular Value Decomposition (GSVD) [5, 23, 60], which provides a simultaneous diagonalisation of a pair of matrices:
@@
The GSVD reverts to the SVD should we can set to recover the SVD of While we assume both in Theorem 3.1 and in the following, analogous results follow for the case Similarly, the GSVD can be stated for singular but this will not be the case for our setting.
@@
It follows from (24) that we have the diagonalisation With this in mind, we call the columns of the left Ritz generalized singular vectors, the columns of the right Ritz generalized singular vectors, and the diagonal entries of the Ritz generalized singular values. Notice that if we take then the Ritz generalized singular values simply are the reciprocals of the Ritz values of Unlike recycling using Ritz vectors, where there was only a single choice regarding what vectors could be used as the recycle space, here we could consider: the left Ritz generalized singular vectors, the right Ritz generalized singular vectors, a linear combination of both the left and right Ritz generalized singular vectors, e.g.
Then the recycle space is given by the columns of the selected matrix corresponding to the smallest or largest Ritz generalized singular values. Thus, in practice, we need only compute the partial GSVD of for components [1, 30, 32], which finds matrices with orthogonal columns, diagonal matrices and full rank matrix such that
@@ 4 Alternative stopping criterion for (11)
where the final equality follows by the definition of the pseudoinverse and (26). In the case that the full GSVD is employed and is square, we have equality throughout the approximation of Thus, we approximate (27) with
We remark that to employ (28) we require an efficient way to determine the current residual vector While RMINRES exploits matrix decompositions to iteratively update the residual norm directly and avoid explicit computation of the residual vector, one can utilize said matrix decompositions encountered in RMINRES to efficiently determine with little overhead; see Appendix A.
@@ 5 Numerical experiments
where denotes a convolution. The parameters to be learned are the exponents defining the regularization parameters and the convolution kernels for so that To guarantee existence and uniqueness of solutions to (29), we fix
Since we are interested in the computational aspects of solving a bilevel learning problem, rather than the generalizabiltiy of the determined parameters, we consider a single data pair and an inpainting test problem, where is a MNIST [35] image (so that ) and is obtained by applying a random subsampling matrix to (so that only of the original pixels are retained) and by adding Gaussian white noise of level The ground truth and associated measurement are displayed in Figure 1. We focus on the bilevel learning problem
to learn convolution kernels of size (so that ).
We employ gradient descent with backtracking linesearch, as outlined in Algorithm 2, to solve (30). The lower level is solved via L-BFGS with backtracking linesearch and limited memory, and we terminate when the gradient norm is smaller than We employ (R)MINRES to solve each Hessian system with a maximum of iterations where, unless specified otherwise, we terminate based on the residual norm with a stopping criterion tolerance of We find that early stopping is always achieved by (R)MINRES. We consider a recycle dimension of with this choice supported by one of the experiments which explores how the dimension of the recycle space affects performance. Various stopping tolerances for (R)MINRES annd L-BFGS were considered and the ones selected were such that the number of iterations of MINRES and L-BFGS were small while the determined solution of the bilevel problem achieved a hypergradient norm comparable to a high-accuracy run.
In Figure 1 we display reconstructions obtained by Total Variation (TV) regularization, a common choice of regularizer for imaging applications, as well FoE regularization with filters determined via the numerical solution to (30). For the TV reconstruction, we selected the regularization parameter achieving the smallest upper level cost among 20 logarithmically equispaced values between and . We see that due to the high undersampling, TV gives a poor reconstruction, whereas a learned regularizer provided by FoE recovers a meaningful reconstruction. A breakdown of the computational timings of the bilevel solve is displayed in Figure 2, where we recall that we employ backtracking linesearch to determine the step length for the upper level problem, and so solve more lower level problems than number of linear systems; see also Algorithms 1 and 2. We see that over 27% of computation time is spent solving the sequence of Hessian systems. Thus, Krylov subspace recycling methods may provide a viable approach to reduce the computational cost by making such task more efficient.
@@
Before describing each test in details, we provide some details about the common computational framework employed for our numerical experiments. We numerically solve the bilevel problem utilizing various recycling strategies, and our Python implementation utilizing PyTorch [48] can be found at github.com/s-j-scott/bilevel-recycling. To the best of our knowledge, unlike the partial eigendecomposition and generalized eigendecomposition, there is no readily available general-purpose implementation of the (partial) GSVD in Python. Thus, to employ our proposed recycling strategy, we follow the GSVD construction as in [45] which involves the computation of two SVDs and a QR decomposition. While matrix decompositions can in principle be performed via a function that returns the respective matrix-vector product, the implementations we utilize require explicit access to the full matrix. Thus, we construct full matrix representations of and in PyTorch. We remark that, to compute Ritz vectors, which only concerns an eigendecomposition, we still compute a full matrix representation of , but then perform the partial eigendecomposition in Python via Scipy.
@@
If we were to solve the entire bilevel problem from scratch with each recycling strategy, a new sequence of Hessians would technically be generated for each strategy. Therefore, in order to have a fair comparison of the performance of the different recycling strategies, we require a single sequence of Hessian systems that can be re-solved using each strategy. We construct this sequence by solving the bilevel problem without any recycling, and tolerances outlined as above.
@@
We are also interested in the potential impact the hypergradients determined via recycling may have on the parameters and associated upper level cost. Since we consider solving a fixed sequence of Hessians that arise from solving the bilevel problem without recycling, we implicitly already have a sequence of parameters one for each linear system. To determine the influence of the hypergradient obtained via recycling, for the th linear system and a given recycling strategy, we start with determined from the non-recycling run of the sequence of systems, employ recycling to solve the associated Hessian for calculate the corresponding hypergradient and then use backtracking linesearch to determine the updated parameter and associated upper level cost. We remark that, in general, is specific to the recycling strategy and can be slightly different to
To compact notation, we denote recycling strategies to be of the form VecType-Size(Side) where

### v2

@@ Abstract
Many optimization problems require hyperparameters, i.e., parameters that must be pre-specified in advance of their solution, such as regularization parameters and parametric regularizers in variational regularization methods for inverse problems, and dictionaries in compressed sensing. A data-driven approach to determine appropriate hyperparameter values is via a nested optimization framework known as bilevel learning. Even when it is possible to employ a gradient-based solver to the bilevel optimization problem, construction of the gradients, known as hypergradients, is computationally challenging, each one requiring both a solution of a minimization problem and a linear system solve. These systems do not change much during the iterations, which motivates us to apply recycling Krylov subspace methods, wherein information from one linear system solve is re-used to solve the next linear system. Existing recycling strategies often employ eigenvector approximations called Ritz vectors. In this work we propose a novel recycling strategy based on a new concept, Ritz generalized singular vectors, which acknowledge the bilevel setting. Additionally, while existing iterative methods primarily terminate according to the residual norm, the newly introduced Ritz generalized singular vectors allow us to define a new stopping criterion that directly approximates the error of the associated hypergradient. The proposed approach is validated through extensive numerical testing in the context of inverse problems in imaging.
@@ 1 Introduction
Large-scale optimization problems that arise in applications such as machine learning [25, 45, 74, 76] and inverse problems [5, 23, 32, 68] often require the pre-specification of certain parameters, such as dictionaries for compressed sensing [15, 44, 48] and sample weights for a biased dataset [63, 69]. Referred to as hyperparameters, it is crucial that appropriate values of these parameters are determined in order to ensure both convergence of the optimization process and that a meaningful solution is obtained. For example, for inverse problems, which arise in many important science and engineering applications such as biomedical, astronomical, and seismic imaging [3, 5, 23, 33], one is tasked to recover some quantity of interest given and indirect measurements thereof. A common approach to determine a meaningful approximation to is to solve an optimization problem such as
where the regularizer is used to encourage a-priori information of the solution in the reconstruction, such as sparsity in some signal domain; see [5, 23, 32, 68]. Here the regularizer is parametrized by some vectors which may represent, for example: the weights for a sum of different regularizers [17]; convolution filters [11, 64]; parameters of a neural network [2, 30, 31, 46]. Determining appropriate values of is crucial to attain a meaningful solution, as a poor choice can lead to a reconstruction that is dominated by noise or over-smoothed [32].
@@
i.e., a linear system whose coefficient matrix is the lower level Hessian. Thus, a gradient-based method applied to (2) involves the solution of a sequence of linear systems of the form (3), which stabilize as the iterations for solving (2) proceed. In applications such as inverse problems, each linear system (3) is typically large scale and thus solved by iterative methods such as Krylov subspace methods. Beyond warm-starting, a way to exploit similarity in a sequence of linear systems is to employ recycling Krylov subspace methods [56, 72, 75], whereby the solution space of one linear system is ‘recycled’ in the solve of the next linear system.
The remaining part of this paper is organized as follows. In Section 2 we provide some background material on gradient-based methods for the solution of bilevel optimization problems, as well as recycling Krylov methods for a sequence of linear systems. In Section 3 we first briefly summarize existing strategies to determine the subspace to be recycled, which are primarily based on Ritz vectors, and then introduce Ritz generalized singular vectors that lead to a novel recycling strategy acknowledging that, in the bilevel optimization setting, solutions of the linear systems are employed in the construction of hypergradients. In Section 4 we show how our proposed recycling strategy allows for a stopping criterion based on approximations of the hypergradient error, as opposed to existing stopping criteria based on the Hessian system residual in some norm [27, 60, 75]. Section 5 reports the results of extensive numerical experiments performed on a couple of test problems, where we show the benefit of using recycling Krylov methods, and compare existing and our new recycling strategy and stopping criterion. Section 6 presents some concluding remarks.
@@ 2.1 Gradient-based solvers for bilevel optimization
where is the Jacobian of with respect to We typically select an upper level loss function such that is easy to compute, e.g., yields An issue with (5) is the application of the Jacobian Indeed, even with existence and uniqueness of we often lack an analytic solution for which we can compute the corresponding Jacobian. However, by the assumptions on an alternative form of the Jacobian motivated by the implicit function theorem [14, 18, 22, 79] can be employed. Namely,
@@
Regarding the computational feasibility of (7), for applications such as imaging problems (whereby is a vectorized image) the Hessian is very large and, in general, lacks any structure. Indeed, we may only have access to a function that implicitly applies the action of the Hessian onto a vector. For these reasons the Hessian is not explicitly inverted in practice, and instead an iterative method is employed to compute the solution of the linear system
@@
At iteration of a gradient-based solver for (4), having the current solution and hypergradient (7) evaluated in (acting as a search direction ), an appropriate steplength can be determined via backtracking linesearch with the Armijo stopping rule. Namely, we find the smallest such that
@@
The caveat of applying the solver described in the previous subsection is that, in order to form each hypergradient (9), one must determine the solution to both a sequence of lower level problems (4b) and Hessian system (8). However, as the iterations of the gradient-based optimization scheme proceed, one would expect the computed hyperparameters to stabilise; in particular the associated sequences of solutions to the lower level and Hessian systems would also stabilise.
We focus our attention on the sequence of linear systems that are solved. To simplify the notation with respect to (8), we denote the Hessian system encountered at the th iteration of a gradient-based method for solving (4), that is, line 4 of Algorithm (2), as
@@ 2.2.1 MINRES
We now focus on how a given linear system in (11) can be solved. To further simplify the notation, we suppress the superscript in (11) and consider the linear system with initial guess One way to solve the linear system is via a Krylov subspace method [29, 65], such as MINRES [54], which builds a Krylov subspace where is the initial residual, and computes a solution that is optimal over
@@
where is the first canonical basis vector. If the residual norm is sufficiently small the algorithm can be terminated, otherwise a new basis vector can be efficiently constructed by the 3-term recurrence (12) and the associated residual norm determined. In particular, MINRES reduces the solution of the original linear system to the solution of a linear least squares problem with tridiagonal coefficient matrix of order , which can be efficiently performed; indeed, the Lanczos relation (12) allows the QR decomposition of to be efficiently updated to yield the QR decomposition of ; see [54] for details.
@@ 2.2.2 RMINRES
RMINRES [75] is a recycling variant of MINRES that finds a correction term of the form
@@
where is a pre-specified recycle space and the columns of form an orthonormal basis of a Krylov subspace which is iteratively constructed. We remind ourselves that, although we suppress the notation in this subsection, in the framework of this paper every variable in (14) - including the recycle space - depends on the iteration of the gradient-based algorithm employed to solve the bilevel problem.
@@
where can be shown to be tridiagonal [72, Section 5.2]. This precise choice of orthogonalisation allows (15) to be decoupled into
@@ 3.1 Existing selection criteria
Existing works that employ recycling strategies are primarily motivated by two settings: a sequence of slowly changing linear systems are solved; or a single linear system is solved but, due to computational memory restrictions, the numerical solver must be restarted mid-solve [51]. As we are interested in solving (11), we tailor discussion to the former. With this in mind, we interpret as the Hessian of the system we are about to solve, i.e., we include a discussion on the implications of this choice later in the section.
Although the two settings are different, existing recycling strategies that exploit approximate invariant subspaces are motivated by approximating the eigendecomposition
where is orthonormal and is a diagonal matrix of the eigenvalues. Indeed, if one has access to the full eigendecomposition (19) then the correction term is directly Due to the large-scale nature of the system and, in general, lack of structure of the Hessian, it is not feasible to compute the eigendecomposition for every Hessian system, hence an iterative method is employed. Even so, since is in the column space of choosing to mimic in some capacity is a natural choice.
@@
and consider a subspace in the range of The columns of are called Ritz vectors, and the associated eigenvalues (entires of the diagonal matrix ) are called the Ritz values. Notice that should we recover the full eigendecomposition Ritz values tend to approximate the extremal eigenvalues of well, but can give poor approximations to interior eigenvalues [52]. For example, eigenvalues can be written as a min-max characterisation of the Rayleigh quotient [57, Theorem 10.2.1] and one of the Ritz values is the largest possible Rayleigh quotient over the given subspace [57, Theorem 11.4.1]. Other characterizations of how Ritz vectors/values can be considered optimal approximations of eigenvectors/values over the subspace spanned by are given in [26, Chapter 8].
It turns out that the Lanczos algorithm for generating Krylov basis vectors is the Rayleigh-Ritz procedure applied to a Krylov subspace [57, Chapter 13], and so the generated Krylov subspace inherently approximates exterior eigenvectors. For this reason, recycling approximations of interior eigenvectors can propagate information least readily available from the Lanczos algorithm. An approximation of interior eigenvalues are Harmonic Ritz values, which are the reciprocal of (ordinary) Ritz values of with respect to Harmonic Ritz values are named as such due to the interpretation of them as a weighted harmonic mean of ’s eigenvalues, in contrast to Ritz values which can be viewed as a weighted arithmetic mean of ’s eigenvalues; see [53] for details. In practice, harmonic Ritz vectors are not determined by forming and computing the associated Ritz vectors, but instead an equivalent generalized eigenvalue problem
@@
Besides recycling (harmonic) Ritz vectors, other recycling strategies are possible. For example, the authors of [40] adopt recycle spaces obtained by combining subspaces of past approximate solutions with approximate invariant subspaces. A recycling strategy inspired by goal-oriented proper orthogonal decomposition from model reduction is proposed in [9]. The authors of [16] use a recycle space defined by approximate right singular vectors of the coefficient matrix. There are also a number of recycling or subspace augmentation strategies explored in the setting of inverse problems, where the main principle underlying the latter is to enrich the approximation subspace for the solution by vectors that may enhance the quality of the reconstruction. For instance, in [39, 58] the recycle space is selected according to specific compression techniques applied to either the projected coefficient matrix or the approximation subspace for the solution; among the latter, choosing the columns of that correspond to the coefficients in of largest magnitude allows to retain the most important columns of that represent the solution. An effective augmented GMRES version was first proposed in [4] and further investigated, among other solvers and even in a continuous setting, in [61]. Similarly, there is an extensive literature about the so-called generalized Krylov subspace methods, where the standard Krylov subspace for the solution (associated to GMRES or CGLS) either appears alongside additional vectors (like in [62]) or is modified by multiplication with iteration-dependent matrices (like in [8]), which are linked to Tikhonov-regularized problems.
@@ 3.2 Acknowledging the bilevel setting
where In practice we first approximate and then perform a premultiplication by The recycling strategies outlined in Section 3.1 use vectors related to projected approximations of and so are oblivious to the premultiplication applied to the determined solution. To acknowledge the bilevel setting, we propose to approximate over the column-space of a full-rank matrix by using the Generalized Singular Value Decomposition (GSVD) [6, 26, 73], which provides a simultaneous as defined in the following statement.
@@
The GSVD reverts to the SVD should , since we can set to recover i.e., the SVD of While we assume both in Theorem 3.1 and in the following, analogous results follow for the case Similarly, the GSVD can be stated for singular but this will not be the case for our setting.
@@
It follows from (24) that we have the diagonalisation With this in mind, we call the columns of the left Ritz generalized singular vectors, the columns of the right Ritz generalized singular vectors, and the diagonal entries of the Ritz generalized singular values. Notice that, if we take then the Ritz generalized singular values simply are the reciprocals of the Ritz values of Unlike recycling using Ritz vectors, where there was only a single choice regarding what vectors could be used as the recycle space, here we could consider: the left Ritz generalized singular vectors, the right Ritz generalized singular vectors, a linear combination of both the left and right Ritz generalized singular vectors, e.g.
Then the recycle space is given by the columns of the selected matrix corresponding to among the smallest and/or largest Ritz generalized singular values. Thus, in practice, we need only compute the partial GSVD of for components [1, 36, 38], which finds matrices with orthogonal columns, diagonal matrices and full rank matrix such that
We conclude this section by emphasising again that building the recycle space upon the newly-introduced Ritz generalized singular vectors and values provides a heuristic way of simultaneously evaluating the action of on the subspace , which contains the approximate solution to the current Hessian system. Although not considered in this paper, there may be other strategies to assess the impact of premultiplying by . For instance, following the optimal truncation results presented for a (sequence of) linear system(s) in [19, 56], it may be possible to prove that retaining the -dimensional subspace of that better aligns (in terms of principal angles between subspaces [26]) with leads to an optimal approximation of within any -dimensional subspace of . In practice, assuming that computing the SVD of is feasible, and denoting by its right singular vector matrix, one may choose , where is the rank- truncated SVD of . Ways of achieving this in a matrix-free setting will be the focus of future research.
@@ 4 Alternative stopping criterion for (11)
where the final equality follows by the definition of the pseudoinverse and (26). In the case that the full GSVD is employed and is square, we have equality throughout the above approximation of Thus, we approximate (27) with
We want to emphasise that, although Ritz vectors can be connected to projected eigenvectors, we have not shown that an analogous statement can be made for Ritz generalized singular vectors and generalized singular vectors. Furthermore, we cannot conclude that the hypergradient error approximation (28) is a projection of (27).
We remark that to employ (28) we require an efficient way to determine the current residual vector While RMINRES exploits matrix decompositions to iteratively update the residual norm directly and avoid explicit computation of the residual vector, one can utilize said matrix decompositions encountered in RMINRES to efficiently determine with little overhead; see supplementary material.
@@ 5 Numerical experiments
where denotes a convolution, and is an elementwise operation (e.g., recovers the squared -norm). The parameters to be learned are the exponents defining the regularization parameters and the filters for so that To guarantee existence and uniqueness of solutions to (29), we fix
We consider training pairs , and focus on the bilevel learning problem
We investigate two applications: an inpainting problem of the MNIST dataset [43], wherein the relatively small dimension of the inverse problem allows for a rigorous empirical exploration; a deconvolution problem of the BSDS300 dataset [49], to illustrate that the proposed recycling strategies remain effective for larger problems.
@@
Before describing each test in detail, we describe the common computational framework employed for our numerical experiments. We numerically solve the bilevel problem utilizing various recycling strategies, and our Python implementation utilizing PyTorch [59] can be found at github.com/s-j-scott/bilevel-recycling. To the best of our knowledge, unlike the partial eigendecomposition and generalized eigendecomposition, there is no readily available general-purpose implementation of the (partial) GSVD in Python. Thus, to employ our proposed recycling strategy, we follow the GSVD construction as in [55], which involves the computation of two SVDs and a QR decomposition. While (partial) matrix decompositions can in principle be performed in a matrix-free fashion (i.e., via a function that returns the respective matrix-vector product), the implementations we utilize require explicit access to the full matrix. Thus, we construct full matrix representations of and in PyTorch. We remark that, to compute Ritz vectors, which only concerns an eigendecomposition, we still compute a full matrix representation of , but then perform the partial eigendecomposition in Python via Scipy.
@@
If we were to solve the entire bilevel problem from scratch with each recycling strategy, a new sequence of Hessians would technically be generated for each strategy. Therefore, in order to have a fair comparison of the performance of the different recycling strategies we require a single sequence of Hessian systems that can be solved using each strategy. To do this, we solve the bilevel problem (30) with , without any recycling and tolerances to be specified.
@@
We are also interested in the potential impact the hypergradients determined via recycling may have on the parameters and associated upper level cost. Since we already solve a fixed sequence of Hessians associated to the bilevel problem without recycling, we implicitly already have a sequence of parameters one for each linear system. To determine the influence of the hypergradient obtained via recycling, for the th linear system and a given recycling strategy, we start with determined from the non-recycling run of the sequence of systems, employ recycling to solve the associated Hessian for calculate the corresponding hypergradient and then use backtracking linesearch to determine the updated parameter and associated upper level cost. We remark that, in general, is specific to the recycling strategy and can be slightly different to
To simplify the notation, we denote recycling strategies to be of the form VecType-Size(Side) where
@@
5.1 MNIST Inpainting
In order to numerically explore the effectiveness of the proposed method and approximations, we consider a small bilevel problem consisting of a single data pair where is a MNIST [43] image (so that ) and is obtained by applying a random subsampling matrix to (so that only of the original pixels are retained) and by adding Gaussian white noise of level The ground truth and associated measurement are displayed in Figure 1. We solve (30) to learn filters of size (so that ) with
We employ gradient descent with backtracking linesearch, as outlined in Algorithm 2, to solve (30). The lower level is solved via L-BFGS (history size ) with backtracking linesearch and we terminate when the gradient norm is smaller than We employ (R)MINRES to solve each Hessian system with a maximum of iterations where, unless specified otherwise, we terminate based on the residual norm with a stopping criterion tolerance of We find that early stopping is always achieved by (R)MINRES. We consider a recycle dimension of with this choice supported by one of the experiments which explores how the dimension of the recycle space affects performance. Various stopping tolerances for (R)MINRES and L-BFGS were considered and the ones selected were such that the number of iterations of MINRES and L-BFGS were small while the determined solution of the bilevel problem achieved a hypergradient norm comparable to a high-accuracy run.
In Figure 1 we display reconstructions of the image used in (30) obtained by Total Variation (TV) regularization, a common choice of regularizer for imaging applications, as well the FoE regularizer. For the TV reconstruction, we selected the regularization parameter achieving the smallest upper level cost among 20 logarithmically equispaced values between and . We see that due to the high undersampling, TV gives a poor reconstruction whereas, in learning filters specific to recovering that image, a learned regularizer provided by FoE recovers a meaningful reconstruction. A breakdown of the computational timings of the bilevel solve is displayed in Figure 2, where we recall that we employ backtracking linesearch to determine the step length for the upper level problem, and so solve more lower level problems than number of linear systems; see also Algorithms 1 and 2. We see that over 26% of computation time is spent solving the sequence of Hessian systems. Thus, Krylov subspace recycling methods may provide a viable approach to reduce the computational cost by making such task more efficient.